# Forward and Backward State Abstractions for Off-policy Evaluation

### v1

@@
Abstract. Off-policy evaluation (OPE) is crucial for evaluating a target policy’s impact offline before its deployment. However, achieving accurate OPE in large state spaces remains challenging. This paper studies state abstractions – originally designed for policy learning – in the context of OPE. Our contributions are three-fold: (i) We define a set of irrelevance conditions central to learning state abstractions for OPE. (ii) We derive sufficient conditions for achieving irrelevance in Q-functions and marginalized importance sampling ratios, the latter obtained by constructing a time-reversed Markov decision process (MDP) based on the observed MDP. (iii) We propose a novel two-step procedure that sequentially projects the original state space into a smaller space, which substantially simplify the sample complexity of OPE arising from high cardinality.
@@ 1 Introduction
Motivation. Off-policy evaluation (OPE) serves as a crucial tool for assessing the impact of a newly developed policy using a pre-collected historical data before its deployment in high-stake applications, such as healthcare (Murphy et al., 2001), recommendation systems (Chapelle & Li, 2011), education (Mandel et al., 2014), dialog systems (Jiang et al., 2021) and robotics (Levine et al., 2020). A fundamental challenge in OPE is its “off-policy” nature, wherein the target policy to be evaluated differs from the behavior policy that generates the offline data. This distributional shift is particularly pronounced in environments with large state spaces of high cardinality. Theoretically, the minimax rate for estimating the target policy’s Q-function decreases rapidly as the state space dimension increases (Chen & Qi, 2022). Empirically, large state space significantly challenges the performance of state-of-the-art OPE algorithms (Fu et al., 2020; Voloshin et al., 2021).
Although different policies induce different trajectories in the large ground state space, they can produce similar paths when restricted to relevant, lower-dimensional state spaces (Pavse & Hanna, 2023). Consequently, applying OPE to these abstract spaces can significantly mitigate the distributional shift between target and behavior policies, enhancing the accuracy in predicting the target policy’s value. This makes state abstraction, designed to reduce state space cardinality, particularly appealing for OPE. However, despite the extensive literature on studying state abstractions for policy learning (see Section 1.1 for details), it has been hardly explored in the context of OPE.
@@
    Introduction of a set of irrelevance conditions for OPE, accompanied by validations of various OPE methods when applied to abstract state spaces under these conditions.
@@
    Derivation of sufficient conditions for state abstractions to achieve irrevelance in Q-functions and marginalized importance sampling (MIS) ratios. A key ingredient of our proposal lies in constructing a time-reversed Markov decision process (MDP, Puterman, 2014) by swapping the future and past. This effectively yields state abstractions that achieve the irrelevance property.
@@
    Development of a novel two-step procedure to sequentially obtain a smaller state space and reduce the sample complexity of OPE. It is also guaranteed to yield a smaller state space compared to existing single-step abstractions.
1.1 Related work
@@
Off-policy evaluation. OPE aims to estimate the average return of a given target policy, utilizing historical data generated by a possibly different behavior policy (Dudík et al., 2014; Uehara et al., 2022). The majority of methods in the literature can be classified into the following three categories:
@@
    Value-based methods that estimate the target policy’s return by learning either a value function (Sutton et al., 2008; Luckett et al., 2019; Li et al., 2024) or a Q-function (Le et al., 2019; Feng et al., 2020; Hao et al., 2021; Liao et al., 2021; Chen & Qi, 2022; Shi et al., 2022) from the data.
@@
    Importance sampling (IS) methods that adjust the observed rewards using the IS ratio, i.e., the ratio of the target policy over the behavior policy, to address their distributional shift. There are two major types: sequential IS (SIS, Precup, 2000; Thomas et al., 2015; Hanna et al., 2019; Hu & Wager, 2023) which employs a cumulative IS ratio, and marginalized IS (Liu et al., 2018; Nachum et al., 2019; Xie et al., 2019; Dai et al., 2020; Yin & Wang, 2020; Wang et al., 2023) which uses the MIS ratio to mitigate the high variance of the SIS estimator.
@@
    Doubly robust methods or their variants that employ both the IS ratio and the value/reward function to enhance the robustness of OPE (Zhang et al., 2013; Jiang & Li, 2016; Thomas & Brunskill, 2016; Farajtabar et al., 2018; Kallus & Uehara, 2020; Tang et al., 2020; Uehara et al., 2020; Shi et al., 2021; Kallus & Uehara, 2022; Liao et al., 2022; Xie et al., 2023).
@@
State abstraction. State abstraction aims to obtain a parsimonious state representation to simplify the sample complexity of reinforcement learning (RL), while ensuring that the optimal policy restricted to the abstract state space attains comparable values as in the original, ground state space. There is an extensive literature on the theoretical and methodological development of state abstraction, particularly bisimulation — a type of abstractions that preserve the Markov property in the abstracted state (Singh et al., 1994; Dean & Givan, 1997; Givan et al., 2003; Ravindran, 2004; Jong & Stone, 2005; Li et al., 2006; Ferns et al., 2004, 2011; Pathak et al., 2017; Wang et al., 2017; Ha & Schmidhuber, 2018; François-Lavet et al., 2019; Gelada et al., 2019; Castro, 2020; Zhang et al., 2020; Allen et al., 2021; Abel, 2022). In particular, Li et al. (2006) analyzed five irrelevance conditions for optimal policy learning. Unlike the aforementioned works that focus on policy learning, we introduce irrelevance conditions for OPE, and propose abstractions that satisfy these irrelevant properties. Meanwhile, the proposed abstraction for achieving irrelevance for the MIS ratio resembles the Markov state abstraction developed by Allen et al. (2021) in the context of policy learning.
More recently, Pavse & Hanna (2023) made a pioneering attempt to study state abstraction for OPE, proving its benefits in enhancing OPE accuracy. However, they primarily focused on MIS estimators. In contrast, our theoretical analysis applies to a broader range of estimators. Moreover, their abstraction did not achieve MIS-ratio irrelevance, nor did they implement the two-step procedure.
Lastly, state abstraction is also related to variable selection (Tangkaratt et al., 2016; Wang et al., 2017; Zhang & Zhang, 2018; Ma et al., 2023) and representation learning for RL (Abel et al., 2016; Shelhamer et al., 2016; Laskin et al., 2020; Uehara et al., 2021).
@@ 2.1 Data generating process, policy, value and IS ratio
Data. Assume the offline dataset $\mathcal{D}$ comprises multiple trajectories, each containing a sequence of state-action-reward triplets $(S_{t},A_{t},R_{t})_{t\geq 1}$ following a finite MDP, denoted by $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\rho_{0},\gamma\rangle$. Here, $\mathcal{S}$ and $\mathcal{A}$ are the discrete state and action spaces, both with finite cardinalities, $\mathcal{T}$ and $\mathcal{R}$ are the state transition and reward functions, $\rho_{0}$ denotes the initial state distribution, and $\gamma\in(0,1)$ is the discount factor.
The data is generated as follows: (i) At the initial time, the state $S_{1}$ is generated according to $\rho_{0}$; (ii) Subsequently, at each time $t$, the agent finds the environment in a specific state $S_{t}\in\mathcal{S}$ and selects an action $A_{t}\in\mathcal{A}$ according to a behavior policy $b$ such that $\mathbb{P}(A_{t}=a|S_{t})=b(a|S_{t})$; (iii) The environment delivers an immediate reward $R_{t}$ with an expected value of $\mathcal{R}(A_{t},S_{t})$, and transits into the next state $S_{t+1}\stackrel{{\scriptstyle d}}{{\sim}}\mathcal{T}(\bullet\mid A_{t},S_{t})$ according to the transition function $\mathcal{T}$. Notice that both the reward and transition functions rely only on the current state-action pair $(S_{t},A_{t})$, independent of the past data history. This ensures that the data satisfies the Markov assumption.
Policy and value. Let $\pi$ denote a given target policy we wish to evaluate. We use $\mathbb{E}^{\pi}$ and $\mathbb{P}^{\pi}$ to denote the expectation and probability assuming the actions are chosen according to $\pi$ at each time. The regular $\mathbb{E}$ and $\mathbb{P}$ without superscript are taking respect to the behavior policy $b$. Our objective lies in estimating the expected cumulative reward under $\pi$, denoted by $J(\pi)=\mathbb{E}^{\pi}\Big{[}\sum_{t=1}^{+\infty}\gamma^{t-1}R_{t}\Big{]}$ using the offline dataset generated under a different policy $b$. Additionally, denote $V^{\pi}$ and $Q^{\pi}$ as the state value function and state-action value function (better known as the Q-function), namely,
@@
IS ratio. We also introduce the IS ratio $\rho^{\pi}(a,s)=\pi(a|s)/b(a|s)$, which quantifies the discrepancy between the target policy $\pi$ and the behavior policy $b$. Furthermore, let $w^{\pi}(a,s)$ denote the MIS ratio $(1-\gamma)\sum_{t\geq 1}\gamma^{t-1}\mathbb{P}^{\pi}(S_{t}=s,A_{t}=a)/\lim_{t\rightarrow\infty}\mathbb{P}(S_{t}=s,A_{t}=a)$. Here, the numerator represents the discounted visitation probability under the target policy $\pi$, a crucial component in policy-based learning for estimating $\pi^{*}$ (Sutton et al., 1999; Schulman et al., 2015). The denominator corresponds to the limiting state-action distribution under the behavior policy. These ratios are fundamental in constructing IS estimators, as detailed in Methods 2 and 3 of Section 2.3.
@@ 2.2 State abstractions for policy learning
Let $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\rho_{0},\gamma\rangle$ be the ground MDP. A state abstraction $\phi$ is a mapping from the state space $\mathcal{S}$ to certain abstract state space $\mathcal{X}=\{\phi(s):s\in\mathcal{S}\}$. Below, we review some commonly studied definitions of state abstraction designed for learning the optimal policy $\pi^{*}$; see Jiang (2018).
@@
The first condition in (2) corresponds to “reward-irrelevance” whereas the second condition represents “transition-irrelevance”. Consequently, Definition 3 defines a “model-based” abstraction, in contrast to “model-free” abstractions considered in Definitions 1 and 2. Notice that the term $\sum_{s^{\prime}\in\phi^{-1}(x^{\prime})}\mathcal{T}(s^{\prime}|a,s)$ – appearing in the second equation of (2) – represents the probability of transitioning to $\phi(S^{\prime})=x^{\prime}$ in the abstract state space. Thus, the second condition essentially requires the abstract next state $\phi(S^{\prime})$ to be conditionally independent of $S$ given $A$ and $\phi(S)$. Assuming $S$ can be decomposed into the union of $\phi(S)$ and $\psi(S)$, which represent relevant features and irrelevant features, respectively. The condition implies that the evolution of those relevant features depends solely on themselves, independent of those irrelevant features. This ensures that the transformed data triplets $(\phi(S),A,R)$ remains an MDP. Meanwhile, the evolution of those irrelevant features may still depend on the relevant features; see Figure 1(a) for an illustration.
It is also known that model-irrelevance implies $Q^{*}$-irrelevance, which in turn implies $\pi^{*}$-irrelevance; see e.g., Theorem 2 in Li et al. (2006). Given that the transformed data remains an MDP under model-irrelevance, one can apply existing state-of-the-art RL algorithms to the abstract state space instead of the original ground space, leading to more effective learning of the optimal policy.
Figure 1: Illustrations of (a) model-irrelevance and (b) backward-model-irrelevance. $\rho_{t}$ is a shorthand for $\rho^{\pi}(A_{t},S_{t})$ for any $t\geq 1$.
@@ 2.3 OPE methodologies
We focus on four OPE methods, covering the three families of estimators introduced in Section 1.1. Each method employs a specific formula to identify $J(\pi)$, which we detail below. The first method is a popular value-based approach – the Q-function-based method. The second and third methods are the two major IS estimators: SIS and MIS. The fourth method is a semi-parametrically efficient doubly robust method, double RL (DRL), known for achieving the smallest possible MSE among a broad class of OPE estimators (Kallus & Uehara, 2020, 2022).
@@
Method 1 (Q-function-based method). For a given Q-function $Q$, define $f_{1}(Q)$ as the estimating function $\sum_{a\in\mathcal{A}}\pi(a|S_{1})Q(a,S_{1})$ with $S_{1}$ being the initial state. By (1) and the definition of $J(\pi)$, it is immediate to see that $J(\pi)=\mathbb{E}[f_{1}(Q^{\pi})]$. This motivates the Q-function-based method which uses a plug-in estimator to approximate $\mathbb{E}[f_{1}(Q^{\pi})]$ and thereby estimates $J(\pi)$. In particular, $Q^{\pi}$ can be estimated by Q-learning type algorithms (e.g., fitted Q-evaluation, FQE, Le et al., 2019), and the expectation can be approximated based on the empirical initial state distribution.
@@
Method 2 (Sequential importance sampling). For a given IS ratio $\rho^{\pi}$, let $\rho_{1:t}^{\pi}$ denote the cumulative IS ratio $\prod_{j=1}^{t}\rho^{\pi}(A_{j},S_{j})$. It follows from the change of measure theorem that the counterfactual reward $\mathbb{E}^{\pi}(R_{t})$ is equivalent to $\mathbb{E}(\rho^{\pi}_{1:t}R_{t})$ whose expectation is taken with respect to the offline data distribution. Assuming all trajectories in $\mathcal{D}$ terminate after a finite time $T$, this allows us to approximate $J(\pi)$ by $\mathbb{E}[f_{2}(\rho^{\pi})]$ where $f_{2}(\rho^{\pi})=\sum_{t=1}^{T}\gamma^{t-1}\rho_{1:t}^{\pi}R_{t}$. The approximation error is bounded by $O(\gamma^{T})$, which decays exponentially fast with respect to $T$. SIS utilizes a plug-in estimator to initially estimate $\rho^{\pi}$ (when the behavior policy is unknown), and subsequently employs this estimator, along with the empirical data distribution, to approximate $\mathbb{E}[f_{2}(\rho^{\pi})]$. However, a notable limitation of this estimator is its rapidly increasing variance due to the use of the cumulative IS ratio $\rho^{\pi}_{1:t}$. Specifically, this variance tends to grow exponentially with respect to $t$, a phenomenon often referred to as the curse of horizon (Liu et al., 2018).
@@
Method 3 (Marginalized importance sampling). The MIS estimator is designed to overcome the limitations of the SIS estimator. It breaks the curse of horizon by incorporating the structure of the MDP model. As noted previously, under the Markov assumption, the reward depends only on the current state-action pair, rather than the entire history. This insight allows us to replace the cumulative IS ratio with the MIS ratio, which depends solely on the current state-action pair. This modification considerably reduces variance because $w^{\pi}$ is no longer history-dependent. Assuming the data trajectory is stationary over time – that is, all state-action-reward $(S,A,R)$ triplets have the same distribution – it can be shown that $J(\pi)=\mathbb{E}[f_{3}(w^{\pi})]$ where $f_{3}(w^{\pi})=(1-\gamma)^{-1}w^{\pi}(A,S)R$ for any triplet $(S,A,R)$. Both $w^{\pi}$ and the expectation can be effectively estimated and approximated using offline data.
@@
Method 4 (Double reinforcement learning). DRL combines Q-function-based method with MIS. Let $f_{4}(Q,w)=f_{1}(Q)+(1-\gamma)^{-1}w(A,S)[R+\gamma\sum_{a}\pi(a|S^{\prime})Q(a,S^{\prime})-Q(A,S)]$, where $f_{1}$ is defined in Method 1 and $(S,A,R,S^{\prime})$ denotes a state-action-reward-next-state tuple. Under the stationarity assumption, it can be shown that $J(\pi)=\mathbb{E}[f_{4}(Q,w)]$ when either $Q=Q^{\pi}$ or $w=w^{\pi}$ (Kallus & Uehara, 2022). DRL proposes to learn both $Q^{\pi}$ and $w^{\pi}$ from the data, employing these estimators to calculate $\mathbb{E}[f_{4}(Q,w)]$ and approximate the expectation with empirical data distribution. The resulting estimator benefits from double robustness: it is consistent when either $Q^{\pi}$ or $w^{\pi}$ is correctly specified.

### v2

# Off-policy Evaluation with Deeply-abstracted States

@@
Abstract. Off-policy evaluation (OPE) is crucial for assessing a target policy’s impact offline before its deployment. However, achieving accurate OPE in large state spaces remains challenging. This paper studies state abstractions – originally designed for policy learning – in the context of OPE. Our contributions are three-fold: (i) We define a set of irrelevance conditions central to learning state abstractions for OPE, and derive a backward-model-irrelevance condition for achieving irrelevance in (marginalized) importance sampling ratios by constructing a time-reversed Markov decision process (MDP). (ii) We propose a novel iterative procedure that sequentially projects the original state space into a smaller space, resulting in a deeply-abstracted state, which substantially simplifies the sample complexity of OPE arising from high cardinality. (iii) We prove the Fisher consistencies of various OPE estimators when applied to our proposed abstract state spaces.
@@ 1 Introduction
Motivation. Off-policy evaluation (OPE) serves as a crucial tool for assessing the impact of a newly developed policy using a pre-collected historical data before its deployment in high-stake applications, such as healthcare (Murphy et al.,, 2001), recommendation systems (Chapelle and Li,, 2011), education (Mandel et al.,, 2014), dialog systems (Jiang et al.,, 2021) and robotics (Levine et al.,, 2020). A fundamental challenge in OPE is its “off-policy” nature, wherein the target policy to be evaluated differs from the behavior policy that generates the offline data. This distributional shift is particularly pronounced in environments with large state spaces of high cardinality. Theoretically, the error bounds for estimating the target policy’s Q-function and value decrease rapidly as the state space dimension increases (Hao et al., 2021a, ; Chen and Qi,, 2022). Empirically, large state space significantly challenges the performance of state-of-the-art OPE algorithms (Fu et al.,, 2020; Voloshin et al.,, 2021).
Although different policies induce different trajectories in the large ground state space, they can produce similar paths when restricted to relevant, lower-dimensional state spaces (Pavse and Hanna,, 2023). Consequently, applying OPE to these abstract spaces can significantly mitigate the distributional shift between target and behavior policies, enhancing the accuracy in predicting the target policy’s value. This makes state abstraction, designed to reduce state space cardinality, particularly appealing for OPE. However, despite the extensive literature on studying state abstractions for policy learning (see Section 2 for details), it has been hardly explored in the context of OPE.
@@
    Introduction of a set of irrelevance conditions for OPE, and derivation of a backward-model-irrelevance condition for state abstractions to achieve irrelevance in marginalized importance sampling ratios by constructing a time-reversed Markov decision process (MDP, Puterman,, 2014) that swaps the future and past.
@@
    Development of a novel iterative procedure to sequentially compress the state space. Specifically, within each iteration, our algorithm consistently produces a state space that is either smaller in size or remains the same. Through its iterative nature, the proposed approach produces a deeply-abstracted state space, which substantially reduces the sample complexity of OPE.
@@
    Validations of various OPE methods when applied to the proposed abstract state spaces.
Organization. The rest of the paper is structured as follows. Section 2 is dedicated to the literature review of related works. MDP-related notions and OPE methodologies relevant to our proposal are recalled in Section 3. Our proposed state abstractions for OPE are presented in Section 4. Section 5 conducts numerical experiments to demonstrate the efficiency of our approach.
2 Related work
@@
Off-policy evaluation. OPE aims to estimate the expected return of a given target policy, utilizing historical data generated by a possibly different behavior policy (Dudík et al.,, 2014; Uehara et al.,, 2022). The majority of methods in the literature can be classified into the following three categories:
@@
    Value-based methods that estimate the target policy’s return by learning either a value function (Sutton et al.,, 2008; Luckett et al.,, 2019; Li et al.,, 2024) or a Q-function (Le et al.,, 2019; Feng et al.,, 2020; Hao et al., 2021b, ; Liao et al.,, 2021; Chen and Qi,, 2022; Shi et al.,, 2022) from the data.
@@
    Importance sampling (IS) methods that adjust the observed rewards using the IS ratio, i.e., the ratio of the target policy over the behavior policy, to address their distributional shift. There are two major types: sequential IS (SIS, Precup,, 2000; Thomas et al.,, 2015; Hanna et al.,, 2019; Hu and Wager,, 2023) which employs a cumulative IS ratio, and marginalized IS (MIS, Liu et al.,, 2018; Nachum et al.,, 2019; Xie et al.,, 2019; Dai et al.,, 2020; Yin and Wang,, 2020; Wang et al.,, 2023) which uses the MIS ratio to mitigate the high variance of the SIS estimator.
@@
    Doubly robust methods or their variants that employ both the IS ratio and the value/reward function to enhance the robustness of OPE (Zhang et al.,, 2013; Jiang and Li,, 2016; Thomas and Brunskill,, 2016; Farajtabar et al.,, 2018; Kallus and Uehara,, 2020; Tang et al.,, 2020; Uehara et al.,, 2020; Shi et al.,, 2021; Kallus and Uehara,, 2022; Liao et al.,, 2022; Xie et al.,, 2023).
@@
State abstraction. State abstraction aims to obtain a parsimonious state representation to simplify the sample complexity of reinforcement learning (RL), while ensuring that the optimal policy restricted to the abstract state space attains comparable values as in the original, ground state space. There is an extensive literature on the theoretical and methodological development of state abstraction, particularly bisimulation — a type of abstractions that preserve the Markov property in the abstracted state (Singh et al.,, 1994; Dean and Givan,, 1997; Givan et al.,, 2003; Ferns et al.,, 2004; Ravindran,, 2004; Jong and Stone,, 2005; Li et al.,, 2006; Ferns et al.,, 2011; Abel et al.,, 2016; Wang et al.,, 2017; Castro,, 2020; Allen et al.,, 2021; Abel,, 2022). In particular, Li et al., (2006) analyzed five irrelevance conditions for optimal policy learning. Unlike the aforementioned works that focus on policy learning, we introduce irrelevance conditions for OPE, and propose abstractions that satisfy these irrelevant properties. Meanwhile, the proposed abstraction for achieving irrelevance for the MIS ratio resembles the Markov state abstraction developed by Allen et al., (2021) in the context of policy learning, while relaxing their requirement for the behavior policy to be Markovian.
More recently, Pavse and Hanna, (2023) made a pioneering attempt to study state abstraction for OPE, proving its benefits in enhancing OPE accuracy. However, they primarily focused on MIS estimators. In contrast, our theoretical analysis applies to a broader range of OPE estimators, covering all three aforementioned categories. Moreover, their abstraction did not achieve MIS-ratio irrelevance. Nor did they implement the iterative procedure.
Lastly, state abstraction is also related to variable selection (Kolter and Ng,, 2009; Geist and Scherrer,, 2011; Geist et al.,, 2012; Nguyen et al.,, 2013; Fan et al.,, 2016; Guo and Brunskill,, 2017; Shi et al.,, 2018; Zhang and Zhang,, 2018; Qi et al.,, 2020; Hao et al., 2021a, ; Ma et al.,, 2023) as well as representation learning for both policy learning (see e.g., Gelada et al.,, 2019; Zhang et al.,, 2020; Uehara et al., 2021b, ) and OPE (see e.g., Wang et al.,, 2021; Chang et al.,, 2022; Ni et al.,, 2023; Pavse and Hanna,, 2024).
@@ 3.1 Data generating process, policy, value and IS ratio
Data. Assume the offline dataset $\mathcal{D}$ comprises multiple trajectories, each containing a sequence of state-action-reward triplets $(S_{t},A_{t},R_{t})_{t\geq 1}$ following a finite MDP, denoted by $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\rho_{0},\gamma\rangle$. Here, $\mathcal{S}$ and $\mathcal{A}$ are the discrete state and action spaces, both with finite cardinalities, $\mathcal{T}$ and $\mathcal{R}$ are the state transition and reward functions, $\rho_{0}$ denotes the initial state distribution, and $\gamma\in(0,1)$ is the discount factor. The data is generated as follows:
  1. 1.
    At the initial time, the state $S_{1}$ is generated according to $\rho_{0}$;
  2. 2.
    Subsequently, at each time $t$, the agent finds the environment in a specific state $S_{t}\in\mathcal{S}$ and selects an action $A_{t}\in\mathcal{A}$ according to a behavior policy $b$ such that $\mathbb{P}(A_{t}=a|S_{t})=b(a|S_{t})$;
  3. 3.
    The environment delivers an immediate reward $R_{t}$ with an expected value of $\mathcal{R}(A_{t},S_{t})$, and transits into the next state $S_{t+1}\stackrel{{\scriptstyle d}}{{\sim}}\mathcal{T}(\bullet\mid A_{t},S_{t})$ according to the transition function $\mathcal{T}$.
Notice that both the reward and transition functions rely only on the current state-action pair $(S_{t},A_{t})$, independent of the past data history. This ensures that the data satisfies the Markov assumption.
Policy and value. Let $\pi$ denote a given target policy we wish to evaluate. We use $\mathbb{E}^{\pi}$ and $\mathbb{P}^{\pi}$ to denote the expectation and probability assuming the actions are chosen according to $\pi$ at each time. The regular $\mathbb{E}$ and $\mathbb{P}$ without superscript are taken with respect to the behavior policy $b$. Our objective lies in estimating the expected cumulative reward under $\pi$, denoted by $J(\pi)=\mathbb{E}^{\pi}\Big{[}\sum_{t=1}^{+\infty}\gamma^{t-1}R_{t}\Big{]}$ using the offline dataset generated under a different policy $b$. Additionally, denote $V^{\pi}$ and $Q^{\pi}$ as the state value function and state-action value function (better known as the Q-function), namely,
@@
IS ratio. We also introduce the IS ratio $\rho^{\pi}(a,s)=\pi(a|s)/b(a|s)$, which quantifies the discrepancy between the target policy $\pi$ and the behavior policy $b$. Furthermore, define the MIS ratio
$w^{\pi}(a,s)=(1-\gamma)\sum_{t\geq 1}\frac{\gamma^{t-1}\mathbb{P}^{\pi}(S_{t}=s,A_{t}=a)}{\displaystyle\lim_{T\rightarrow\infty}\mathbb{P}(S_{T}=s,A_{T}=a)}.$ (2)
Here, the numerator represents the discounted probability of visiting a given state-action pair under the target policy $\pi$, a crucial component in policy-based learning for estimating $\pi^{*}$ (Sutton et al.,, 1999; Schulman et al.,, 2015). The denominator corresponds to the limiting state-action distribution under the behavior policy. These ratios are fundamental in constructing IS estimators, as detailed in Methods 2 and 3 of Section 3.3.
@@ 3.2 State abstractions for policy learning
Let $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\rho_{0},\gamma\rangle$ be the ground MDP. A state abstraction $\phi$ is a mapping from the state space $\mathcal{S}$ to certain abstract state space $\mathcal{X}=\{\phi(s):s\in\mathcal{S}\}$. Below, we review some commonly studied definitions of state abstraction designed for learning the optimal policy $\pi^{*}$; see Jiang, (2018).
@@
The first condition in (3) corresponds to “reward-irrelevance” whereas the second condition represents “transition-irrelevance”. Consequently, Definition 3 defines a “model-based” abstraction, in contrast to “model-free” abstractions considered in Definitions 1 and 2. Notice that the term $\sum_{s^{\prime}\in\phi^{-1}(x^{\prime})}\mathcal{T}(s^{\prime}|a,s)$ – appearing in the second equation of (3) – represents the probability of transitioning to $\phi(S^{\prime})=x^{\prime}$ in the abstract state space. Thus, the second condition essentially requires the abstract next state $\phi(S^{\prime})$ to be conditionally independent of $S$ given $A$ and $\phi(S)$. Assuming $S$ can be decomposed into the union of $\phi(S)$ and $\psi(S)$, which represent relevant features and irrelevant features, respectively. This condition implies that the evolution of those relevant features depends solely on themselves, independent of those irrelevant features. This ensures that the transformed data triplets $(\phi(S),A,R)$ remains an MDP. Meanwhile, the evolution of those irrelevant features may still depend on the relevant features; see Figure 1(a) for an illustration.
It is also known that model-irrelevance implies $Q^{*}$-irrelevance, which in turn implies $\pi^{*}$-irrelevance; see e.g., Theorem 2 in Li et al., (2006). Given that the transformed data remains an MDP under model-irrelevance, one can apply existing state-of-the-art RL algorithms to the abstract state space instead of the original ground space, leading to more effective learning of the optimal policy.
Figure 1: Illustrations of (a) model-irrelevance and (b) backward-model-irrelevance. $S_{t}$ is decomposed into the union of $\phi(S_{t})$ (relevant features) and $\psi(S_{t})$ (irrelevant features).
@@ 3.3 OPE methodologies
We focus on four OPE methods, covering the three families of estimators introduced in Section 2. Each method employs a specific formula to identify $J(\pi)$, which we detail below. The first method is a popular value-based approach – the Q-function-based method. The second and third methods are the two major IS estimators: SIS and MIS. The fourth method is a semi-parametrically efficient doubly robust method, double RL (DRL), known for achieving the smallest possible MSE among a broad class of OPE estimators (Kallus and Uehara,, 2020, 2022; Liao et al.,, 2022).
@@
Method 1 (Q-function-based method). For a given Q-function $Q$, define $f_{1}(Q)$ as the estimating function $\sum_{a\in\mathcal{A}}\pi(a|S_{1})Q(a,S_{1})$ with $S_{1}$ being the initial state. By (1) and the definition of $J(\pi)$, it is immediate to see that $J(\pi)=\mathbb{E}[f_{1}(Q^{\pi})]$. This motivates the Q-function-based method which uses a plug-in estimator to approximate $\mathbb{E}[f_{1}(Q^{\pi})]$ and estimate $J(\pi)$. In particular, $Q^{\pi}$ can be estimated by Q-learning type algorithms (e.g., fitted Q-evaluation, FQE, Le et al.,, 2019), and the expectation can be approximated based on the empirical initial state distribution.
@@
Method 2 (Sequential importance sampling). For a given IS ratio $\rho^{\pi}$, let $\rho_{1:t}^{\pi}$ denote the sequential IS ratio $\prod_{j=1}^{t}\rho^{\pi}(A_{j},S_{j})$. It follows from the change of measure theorem that the counterfactual reward $\mathbb{E}^{\pi}(R_{t})$ is equivalent to $\mathbb{E}(\rho^{\pi}_{1:t}R_{t})$ whose expectation is taken with respect to the offline data distribution. Assuming all trajectories in $\mathcal{D}$ terminate after a finite time $T$, this allows us to represent $J(\pi)$ by $\mathbb{E}[f_{2}(\rho^{\pi})]$ where $f_{2}(\rho^{\pi})=\sum_{t=1}^{T}\gamma^{t-1}\rho_{1:t}^{\pi}R_{t}$. SIS utilizes a plug-in estimator to first estimate $\rho^{\pi}$ (when the behavior policy is unknown), and then employs this estimator, along with the empirical data distribution, to approximate $\mathbb{E}[f_{2}(\rho^{\pi})]$. However, a notable limitation of this estimator is its rapidly increasing variance due to the use of the SIS ratio $\rho^{\pi}_{1:t}$. Specifically, this variance tends to grow exponentially with respect to $t$, a phenomenon often referred to as the curse of horizon (Liu et al.,, 2018).
@@
Method 3 (Marginalized importance sampling). The MIS estimator is designed to overcome the limitations of the SIS estimator. It breaks the curse of horizon by taking the structure of the MDP model into account. As noted previously, under the Markov assumption, the reward depends only on the current state-action pair, rather than the entire history. This insight allows us to replace the SIS ratio with the MIS ratio, which depends solely on the current state-action pair. This modification considerably reduces variance because $w^{\pi}$ is no longer history-dependent. Assuming the data trajectory is stationary over time – that is, all state-action-reward $(S,A,R)$ triplets have the same distribution – it can be shown that $J(\pi)=\mathbb{E}[f_{3}(w^{\pi})]$ where $f_{3}(w^{\pi})=(1-\gamma)^{-1}w^{\pi}(A,S)R$ for any triplet $(S,A,R)$. Both $w^{\pi}$ and the expectation can be effectively approximated using offline data.
@@
Method 4 (Double reinforcement learning). DRL combines the Q-function-based method with MIS. Let $f_{4}(Q,w)=f_{1}(Q)+(1-\gamma)^{-1}w(A,S)[R+\gamma\sum_{a}\pi(a|S^{\prime})Q(a,S^{\prime})-Q(A,S)]$, where $f_{1}$ is defined in Method 1 and $(S,A,R,S^{\prime})$ denotes a state-action-reward-next-state tuple. Under the stationarity assumption, it can be shown that $J(\pi)=\mathbb{E}[f_{4}(Q,w)]$ when either $Q=Q^{\pi}$ or $w=w^{\pi}$ (Kallus and Uehara,, 2022). DRL proposes to learn both $Q^{\pi}$ and $w^{\pi}$ from the data, employing these estimators to calculate $\mathbb{E}[f_{4}(Q,w)]$ and approximate the expectation with empirical data distribution. The resulting estimator benefits from double robustness: it is consistent when either $Q^{\pi}$ or $w^{\pi}$ is correctly specified.